# Video-Guided Foley Sound Generation with Multimodal Controls

### v2

@@ 2 Related Work
Figure 3: Method overview. We train our model jointly on a standard audio-video dataset VGGSound for VT2A generation and a high-quality audio-text dataset Sound ideas for T2A generation. We encode the input audio into latents, adding noise to a portion of them. The silent video is encoded into visual features, concatenated with the audio latents along the channel dimension. The text input, including a quality tag, is encoded through a text encoder and applied via cross-attention.
@@ 3.2 Multimodal Conditioning and Training
To enhance audio generation quality and text control, we train our model jointly on an audio-visual dataset, VGGSound [12], and a professional audio-text dataset sampled from Sound Ideas [^1] https://www.sound-ideas.com/. The audio in these datasets is quite different, since VGGSound is “in-the-wild” and relatively noisy, whereas Sound Ideas contains professionally recorded sound effects and thus is generally higher quality and full bandwidth. To give the model control over audio quality, we assign quality tags to both datasets: VGGSound examples are paired with a “low quality” tag in prompts, e.g., “category, low quality”. In contrast, Sound Ideas examples are labeled with a “high quality” tag in their captions. During training, we randomly drop out the caption or quality tag to allow the model to disentangle audio quality from semantics.
@@
The model is jointly trained for video-text-to-audio generation on VGGSound, containing 168K samples, and for text-to-audio generation on Sound-ideas, with 400K samples. We balance the datasets to ensure robust training and train the model for 600K iterations with a batch size of 128. During training, audio latents are masked with a probability of 0.25, where the masking spans 0 to 2 seconds. We also randomly drop the video, caption, and quality tag conditions during the training. We then finetune this pretrained model for 50k iterations on a curated subset of VGGSound, selected for high audio-visual correspondence using an ImageBind [32] score threshold of 0.3. During inference, we use the DDIM [84] method and apply a classifier-free guidance scale ranging from 3.0 to 5.0 to denoise over 100 steps. See Appendix A.1 for more details.
@@ Appendix A.1 Implementation Details
We use the AdamW optimizer [47, 61] with a learning rate of $10^{-4}$ and apply a cosine decay schedule. Training begins with a linear warm-up phase for the first 4K iterations, followed by 599.6K iterations. We train our model with Exponential Moving Average (EMA) [67] with EMA decay of 0.99. Throughout the training, we randomly sample from combined datasets where 60% of training examples are from VGGSound and 40% from Sound Ideas. Within VGGSound samples, 60% are dedicated to video-text-tag-to-audio generation, the rest 40% are evenly distributed across different dropout variants (i.e., video+tag, video+text, video-only, text+tag, text-only, tag-only, unconditional). For Sound Ideas samples, 60% are allocated to text-tag-to-audio generation, with the remaining cases divided as follows: 10% for text-only, 15% for tag-only, and 15% for unconditional audio generation.

### v3

@@ 2 Related Work
Figure 3: Method overview. We train our model jointly on a standard audio-video dataset VGGSound for VT2A generation and a high-quality audio-text dataset HQ-SFX for T2A generation. We encode the input audio into latents, adding noise to a portion of them. The silent video is encoded into visual features, concatenated with the audio latents along the channel dimension. The text input, including a quality tag, is encoded through a text encoder and applied via cross-attention.
@@ 3.2 Multimodal Conditioning and Training
To enhance audio generation quality and text control, we train our model jointly on an audio-visual dataset, VGGSound [12], and a large proprietary licensed dataset of high-quality sound effects with text captions, termed as HQ-SFX. The audio in these datasets is quite different, since VGGSound is “in-the-wild” and relatively noisy, whereas HQ-SFX contains professionally recorded sound effects and thus is generally higher quality and full bandwidth. To give the model control over audio quality, we assign quality tags to both datasets: VGGSound examples are paired with a “low quality” tag in prompts, e.g., “category, low quality”. In contrast, HQ-SFX examples are labeled with a “high quality” tag in their captions. During training, we randomly drop out the caption or quality tag to allow the model to disentangle audio quality from semantics.
@@
The model is jointly trained for video-text-to-audio generation on VGGSound, containing 168K samples, and for text-to-audio generation on the HQ-SFX dataset, with 400K samples. We balance the datasets to ensure robust training and train the model for 600K iterations with a batch size of 128. During training, audio latents are masked with a probability of 0.25, where the masking spans 0 to 2 seconds. We also randomly drop the video, caption, and quality tag conditions during the training. We then finetune this pretrained model for 50k iterations on a curated subset of VGGSound, selected for high audio-visual correspondence using an ImageBind [32] score threshold of 0.3. During inference, we use the DDIM [84] method and apply a classifier-free guidance scale ranging from 3.0 to 5.0 to denoise over 100 steps. See Appendix A.1 for more details.
@@ Appendix A.1 Implementation Details
We use the AdamW optimizer [47, 61] with a learning rate of $10^{-4}$ and apply a cosine decay schedule. Training begins with a linear warm-up phase for the first 4K iterations, followed by 599.6K iterations. We train our model with Exponential Moving Average (EMA) [67] with EMA decay of 0.99. Throughout the training, we randomly sample from combined datasets where 60% of training examples are from VGGSound and 40% from HQ-SFX. Within VGGSound samples, 60% are dedicated to video-text-tag-to-audio generation, the rest 40% are evenly distributed across different dropout variants (i.e., video+tag, video+text, video-only, text+tag, text-only, tag-only, unconditional). For HQ-SFX samples, 60% are allocated to text-tag-to-audio generation, with the remaining cases divided as follows: 10% for text-only, 15% for tag-only, and 15% for unconditional audio generation.